OpenStudy (anonymous):

help please :3? y + 1.2y +1.2z A) 2.4yz B)1.3y + 1.2z C) 1.2y2 + 1.2z D) 2.2y + 1.2z thanks :D

OpenStudy (anonymous):

combine link terms

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

y + 1.2y =?

OpenStudy (anonymous):

um 1.2yy ?

OpenStudy (anonymous):

1+1.2=?

OpenStudy (anonymous):

2.2

OpenStudy (anonymous):

so y + 1.2y=?

OpenStudy (anonymous):

I don't know what to do with the y

OpenStudy (anonymous):

its just like you factor a y y(1+1.2)=2.2y so your answer is 2.2y + 1.2z

OpenStudy (anonymous):

oh :o do i have to solve that on to simplify

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

thanks :p

OpenStudy (anonymous):

@surdawi why did y count as one

OpenStudy (anonymous):

1*y=y; you factored a y

OpenStudy (anonymous):

6r + r – 5r 6r +r = 7r 7r - 5r = 3? or 3r

OpenStudy (anonymous):

is that right the origonal problem is 6r + r – 5r

OpenStudy (anonymous):

6r + r – 5r factor an r r*(6 + 1 - 5) what should you get?

OpenStudy (anonymous):

UM 7 -5 = 3r

OpenStudy (anonymous):

2r